OpenStudy (anonymous):

Function Help, 2 more questions

OpenStudy (anonymous):

Let f(x) = x + 8 and g(x) = x2 - 6x - 7. Find f(g(2)). -7 -3 10 33

OpenStudy (anonymous):

f(x)= 12 g(x) =-15 -3 I tried but I dont know if its right

OpenStudy (cwrw238):

first plug in 2 instead of x in the formula for g(x) and work it out then plug in the value you found into f(x)

OpenStudy (cwrw238):

-15 is right for g(2)

OpenStudy (cwrw238):

so know you have -15 + 8

OpenStudy (anonymous):

Ok x2 - 6x - 7 4 - 12 -7 = -15 x+8 -15 +8= -7?

OpenStudy (cwrw238):

f(g(2)) means the value of f(x) where g(2) replaces the x (in f(x))

OpenStudy (cwrw238):

-7 is correct

OpenStudy (anonymous):

thanks, that wasn't too hard :p

OpenStudy (anonymous):

can i post my last question here or make new thread?

OpenStudy (cwrw238):

yw no - once you understand the notation its easy

OpenStudy (cwrw238):

post it

OpenStudy (anonymous):

Let f(x) = -4x + 7 and g(x) = 10x - 6. Find f(g(x)). -40x + 24 -40x + 31 -40x + 64 -40x + 70

OpenStudy (cwrw238):

here you replace the x in f(x) with 10x - 6 and simplify

OpenStudy (cwrw238):

I'll start it off:- = -4(10x - 6) + 7

OpenStudy (anonymous):

(-40x + 24) + 7

OpenStudy (cwrw238):

ok now you can simplify further

OpenStudy (anonymous):

(-40x + 24) + 7 -7 -7 -40x + 17? (-40x + 24) +7 -24 -24 -40x - 17?

OpenStudy (anonymous):

I don't know which one

OpenStudy (cwrw238):

there was no need to write in the brackets after you did the multiplication it is just -40x + 24 + 7 = -40x + 31

OpenStudy (anonymous):

or is it +7 to 24 to get -40x + 31

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

didn't think first answer sounded right when i did it, thanks again

OpenStudy (cwrw238):

after you distribute the -4 over the brackets , the brackets disappear

OpenStudy (anonymous):

okay. Let f(x) = twenty-four over the quantity of 3 x minus two. Find f(-2). -6 -3 <--? 3 6 is this one right since you're here

OpenStudy (cwrw238):

is it